A mother wants to invest $11,000.00 for her son's future education. She invests a portion of the money in a bank certificate of deposit (CD account) which earns 4% and the remainder in a savings bond that earns 7%. If the total interest earned after one year is $660.00. How much money was invested in the CD account
Answer by mananth(16946)

CD 4.00% per annum ------------- Amount invested =x
Savings bond 7.00% per annum ------------ Amount invested = y
11000
Interest----- 660

CD 4.00% per annum ---x
Savings bond 7.00% per annum ---y
Sum of investments
x + y= 11000 -------------1
Interest earned on both investments
4.00% x + 7.00% y= 660
Multiply by 100
4 x + 7 y= 66000.00 --------2
Multiply (1) by -4
we get
-4 x -4 y= -44000.00
Add this to (2)
0 x 3 y= 22000
divide by 3
y = 7333.33
CD 7.00% $ 7333.33
Savings bond 4.00% $ 3666.67

CHECK
7333.33 --------- 7.00% ------- 513.33
3666.67 ------------- 4.00% ------- 146.67
Total -------------------- 660.00

m.ananth@hotmail.ca